You are an industry analyst that specializes in an industry where the market inverse demand is P = 100 - 2Q. The external marginal cost of producing the product is MCExternal = 8Q, and the internal cost is MCInternal = 18Q.

Instructions: Enter your responses rounded to the nearest two decimal places.

a. What is the socially efficient level of output? _____units
b. Given these costs and market demand, how much output would a competitive industry produce? ____units
c. Given these costs and market demand, how much output would a monopolist produce?_____ units
d. Which of the following are actions the government could take to induce firms in this industry to produce the socially efficient level of output.
Answer #1

A) Setting P= MC internal + MC external we have,

100-2Q= 8Q+18Q

Q= 100/28 = 3.57

B) competitive industry will produce

P= MC internal

100-2Q=18Q

Q= 5

C) For monopolist

MR=MC

MR = 100- 4Q

MC =18 Q

Therefore, 100= 22Q

Q= 4.54